OpenStudy (anonymous):

Rectangle R has varying length l and width w but a constant area f 4ft^2. A. Express the perimeter P as a function of l. What kind of function is P? What is its domain? B. Describe the asymptotic behavior of P. What can you say about R because of the behavior?

OpenStudy (anonymous):

34

OpenStudy (anonymous):

Can you explain?

OpenStudy (anonymous):

You use a calculator

OpenStudy (anonymous):

how'd you get 34?

OpenStudy (anonymous):

i guessed

OpenStudy (anonymous):

@360REVOLUTION @Isaiah.Feynman @willreel @133 help please?any idea

OpenStudy (anonymous):

@ganeshie8 could you help me?

OpenStudy (lastdaywork):

Let area = A = l * w P = 2l + 2w = 2l + 2A/l

OpenStudy (anonymous):

ooookkaaayyy...then what?

OpenStudy (lastdaywork):

I dont know what to call it - P = 2l + 2A/l First term is linear and second is (rectangular) hyperbolic.

OpenStudy (lastdaywork):

I think you should try to find its domain yourself..

OpenStudy (anonymous):

ya ok. could you help with part b?

OpenStudy (lastdaywork):

Do you know the definition of asymptote ??

OpenStudy (anonymous):

it doesn't meet any distance and it keeps a continuous curve?

OpenStudy (lastdaywork):

See the graph and answer (b) http://www.wolframalpha.com/input/?i=y%3D2x+%2B+8%2Fx

OpenStudy (anonymous):

what would I say about the rectangle because of this?